# myStyleroundcorner=5pt,backgroundcolor=blue!75!cyan!05,linecolor=blue!40!black,linewidth=1.5pt

### v1

@@ Abstract
We consider the problem of a graph subjected to adversarial perturbations, such as those arising from cyber-attacks, where edges are covertly added or removed. The adversarial perturbations occur during the transmission of the graph between a sender and a receiver. To counteract potential perturbations, we explore a repetition coding scheme with sender-assigned binary noise and majority voting on the receiver’s end to rectify the graph’s structure. Our approach operates without prior knowledge of the attack’s characteristics. We provide an analytical derivation of a bound on the number of repetitions needed to satisfy probabilistic constraints on the quality of the reconstructed graph. We show that the method can accurately decode graphs that were subjected to non-random edge removal, namely, those connected to vertices with the highest eigenvector centrality, in addition to random addition and removal of edges by the attacker.
@@ I Introduction
Machine learning techniques that use deep neural networks, notably graph neural networks (GNNs), have received great attention in various areas of physics (Xie and Grossman, 2018; Huang and Moore, 2021; Desai et al., 2021; Soltani et al., 2022; Ehrke et al., 2023; Yu et al., 2024). The interest stems from graphs encoding complex relationships between the constituents of the inputs that traditional vector/matrix representations (resulting in simple Euclidean relationships between the constituents) cannot. This work addresses the problem of correcting adversarial perturbations to graphs, a topic of great importance in defending against cyber-attacks (Vivek et al., 2019). Specifically, we consider an adversarial environment that includes an attacker and a defender. The defender is a GNN that takes graphs as inputs. The attacker can manipulate the inputs (graphs) of the GNNs by adding or removing edges Dai et al. (2018); Zügner et al. (2018); Bojchevski and Günnemann (2019); Wu et al. (2019); Ma et al. (2020). The purpose of such attacks is to trigger misclassification in the GNNs.
Specifically, we provide a mathematical analysis of error-correcting codes used to undo the attack (or de-noise the input graphs). We consider a scenario in which the defender involves two parties, a sender and receiver, where in between, the graph is susceptible to an adversarial attack. We employ an encoding scheme in which the sender randomly perturbs $K$ copies of the graph, and the receiver applies a majority voting scheme on the $K$ copies received (after being subjected to adversarial perturbations); see Fig. 1 for illustration.
@@ II.2 $\rho$-robustness
In our setting with graph data, the natural choice for $p$ is 1, i.e., the norm is the $\ell_{1}$-norm, and the perturbation uses the XOR operator $\oplus$ [^1] Recall that $0\oplus 0=0$, $0\oplus 1=1$, $1\oplus 0=1$, $1\oplus 1=0$. to append the attack vector rather than simple addition.
@@
where $w_{\mathcal{N}(0,\sigma^{2}\mathbf{I})}(\cdot)$ is the probability density function (PDF) of a zero-mean Gaussian random vector with covariance matrix $\sigma^{2}\mathbf{I}$ and $\langle\cdot\rangle_{w}$ denotes mathematical expectation with respect to the PDF $w$. We will omit the subscript $w$ where the underlying probability law is understood from context. Suppose $y^{*}$ is the class with the highest probability predicted by $g$ when the input is $\mathbf{s}$, i.e.,
@@
for all perturbations $\xi$ that satisfy $\|\xi\|_{2}\leq\rho$ with certification radius given by
@@
[style = myStyle]
Typically, $\rho$ is very small. Larger values of $\rho$ render the classifiers prone to misclassification, albeit robustly, due to the smoothing of the classification boundaries in the input space. Our approach complements $\rho$-robustness by ensuring input corrections, allowing for training with small values of $\rho$, thereby mitigating misclassification due to adversarial training.
@@ III.1 The encoding scheme (sender)
Our encoding scheme assumes that $\nu$, $N$, and $K$ are known to the defender but hidden from the attacker. The perturbation probability $\beta$ is not known to the defender. The attacker can see the size of the perturbed signal $NK$ but not the factors $N$ and $K$.
@@ V Experiments
To test the method, we consider an attacker performing two types of perturbations to a randomly generated graph: the first is a random perturbation to the edges of the graph, each with probability 0.2, and the second is one where the attacker disconnects the vertex with the highest eigenvector centrality. Throughout our experiments, we treat $\rho=0.05$ as a reasonable robustness target; that is, the classifiers can handle errors of up to 5% in the inputs.
@@
We summarize our results using empirical probabilities, namely kernel density estimators (KDEs):
@@
We explore this further for small graphs below but first quantify (more precisely) $\mathbb{P}(\frac{1}{N}\|\widehat{\mathbf{s}}_{K}-\mathbf{s}\|_{1}\leq\rho)$ through our simulations and $\widehat{F}_{M}$.
@@
The final experiments test the limits of the proposed approach, namely for small graphs with $N=50$. For $N=50$, we observe a second error mode in the empirical PDF in Fig. 8a when $\nu=0.1$. The error mode persists across all values of $K$. This error mode corresponds to edges associated with the vertex with the highest eigenvector centrality, which does not sufficiently change across the $K$ repetitions. Setting $\nu=0.25$, we see in Fig. 8b that this second error mode vanishes. See Fig. 9 for the ECDF.
@@
From Fig. 9, our theory’s conservative estimate of $K$ aligns with the required repetitions needed to achieve a robustness close to (but smaller than) $\rho=0.05$. Thus, the theory produces estimates that are geared towards more aggressive attacks.
@@ VI Conclusion
For attacks that target specific edges (e.g., Xu et al. (2019)), the defender-added perturbation $\vec{\zeta}$ becomes necessary. It serves the role of concealing the graph structure from potential attackers. The bound on the number of repetitions derived above tends to be robust to more aggressive attacks. This is understandable as it aims to be robust to attacks of unknown nature. An interesting follow-up study could investigate embedding encoding-decoding layers in the GNN classifiers at training time to decode the inputs before classification.

### v2

@@ Abstract
We consider the problem of a graph subjected to adversarial perturbations, such as those arising from cyber-attacks, where edges are covertly added or removed. The adversarial perturbations occur during the transmission of the graph between a sender and a receiver. To counteract potential perturbations, this study explores a repetition coding scheme with sender-assigned noise and majority voting on the receiver’s end to rectify the graph’s structure. The approach operates without prior knowledge of the attack’s characteristics. We analytically derive a bound on the number of repetitions needed to satisfy probabilistic constraints on the quality of the reconstructed graph. The method can accurately and effectively decode Erdős-Rényi graphs that were subjected to non-random edge removal, namely, those connected to vertices with the highest eigenvector centrality, in addition to random addition and removal of edges by the attacker. The method is also effective against attacks on large scale-free graphs generated using the Barabási-Albert model but require a larger number of repetitions than needed to correct Erdős-Rényi graphs.
@@ I Introduction
Machine learning techniques that use deep neural networks, notably graph neural networks (GNNs), have received great attention in various areas of physics (Xie and Grossman, 2018; Huang and Moore, 2021; Desai et al., 2021; Soltani et al., 2022; Ehrke et al., 2023; Yu et al., 2024). The interest stems from graphs encoding complex relationships between the constituents of the inputs that traditional vector/matrix representations (resulting in simple Euclidean relationships between the constituents) cannot. This work addresses the problem of correcting adversarial perturbations to graphs, a topic of great importance in defending against cyber-attacks (Vivek et al., 2019). Specifically, we consider an adversarial environment that includes an attacker and a defender. The defender is a GNN that takes graphs as inputs. The attacker can manipulate the inputs (graphs) of the GNNs by adding or removing edges Zügner et al. (2018); Dai et al. (2018); Bojchevski and Günnemann (2019); Wu et al. (2019); Ma et al. (2020). The purpose of such attacks is to trigger misclassification in the GNNs. Studies Zügner et al. (2018); Dai et al. (2018); Bojchevski and Günnemann (2019); Wu et al. (2019); Ma et al. (2020) focused on the attack models and techniques to solve them. Wu et al. (2019) and Zhang and Zitnik (2020) developed adversarial training methods that defend GNNs against poisoned inputs during training. Gosch et al. (2024) derived theoretical guarantees of robustness for adversarial training methods in prior works. This paper aims to complement these works on adversarial training.
Specifically, this paper provides a mathematical analysis of error-correcting codes used to undo the attack (or de-noise the input graphs). We consider a scenario in which the defender involves two parties, a sender and receiver, where in between, the graph is susceptible to an adversarial attack. The method employs an encoding scheme in which the sender randomly perturbs $K$ copies of the graph, and the receiver applies a majority voting scheme on the $K$ copies received (after being subjected to adversarial perturbations); see Fig. 1 for illustration.
@@ II.2 $\rho$-robustness
In our setting with graph data, the natural choice for $p$ is 1, i.e., the norm is the $\ell_{1}$-norm, and the perturbation uses the XOR operator $\oplus$ to append the attack vector.
@@
where $w_{\mathcal{N}(0,\sigma^{2}\mathbf{I})}(\cdot)$ is the probability density function (PDF) of a zero-mean Gaussian random vector with covariance matrix $\sigma^{2}\mathbf{I}$ and $\langle\cdot\rangle_{w}$ denotes mathematical expectation with respect to the PDF $w$. We will omit the subscript $w$ where the underlying probability law is understood from context. Suppose $y^{*}$ is the class with the highest probability predicted by $g$ for input $\mathbf{s}$, i.e.,
@@
for all perturbations $\xi$ that satisfy $\|\xi\|_{2}\leq\rho$ with certification radius
@@
[style = myStyle] Typically, $\rho$ is small. Increasing $\rho$ increases the risk of misclassification, albeit robustly, due to smoothing of the classification boundaries in the input space. Our approach complements $\rho$-robustness via input corrections, allowing for small values of $\rho$, mitigating misclassification due to smoothing.
@@ III.1 The encoding scheme (sender)
Our encoding scheme assumes that $\nu$, $N$, and $K$ are known to the defender but hidden from the attacker. The perturbation probability $\beta$ is not known to the defender. The attacker can see the size of the perturbed signal $NK$ but not the factors $N$ and $K$. The time complexity of the encoding scheme is $O(NK)$ since generating a Bernoulli random variate has the complexity of generating a uniform pseudo-random number on $(0,1)$, which is $O(1)$. As $N\propto|\mathcal{V}|^{2}$, the time complexity of the encoding scheme scales with the graph size as $O(K|\mathcal{V}|^{2})$. However, the encoding scheme can be parallelized. One can achieve a best-case time complexity of $O(K)$ by encoding each of the $N$ blocks in parallel.
@@ III.2 The decoding scheme (receiver)
Since the entries of $\mathbf{r}$ are binary, the worst-case time complexity of the majority vote for each $n$ is $O(K)$. Thus, like the encoding scheme, decoding scales with the graph size as $O(K|\mathcal{V}|^{2})$. However, unlike encoding, decoding cannot be parallelized.
@@ V Experiments
To test the method, we consider an attacker performing two types of perturbations to a randomly generated graph, i.e., the Erdős-Rényi model: the first is a random perturbation to the edges of the graph, each with probability 0.2, and the second is one where the attacker disconnects the vertex with the highest eigenvector centrality. Throughout our experiments, we treat $\rho=0.05$ as a reasonable robustness target; that is, the classifiers can handle errors of up to 5% in the inputs.
@@
We summarize the results using empirical probabilities, namely kernel density estimators (KDEs):
@@
We’ll explore this further for small graphs below but let’s first quantify $\mathbb{P}(\frac{1}{N}\|\widehat{\mathbf{s}}_{K}-\mathbf{s}\|_{1}\leq\rho)$ more precisely through the simulations and $\widehat{F}_{M}$.
@@
For $N=50$, we observe a second error mode in the empirical PDF in Fig. 8a when $\nu=0.1$. The error mode persists across all values of $K$. This error mode corresponds to edges associated with the vertex with the highest eigenvector centrality, which does not sufficiently change across the $K$ repetitions. Setting $\nu=0.25$, we see in Fig. 8b that this second error mode vanishes. See Fig. 9 for the ECDF. From Fig. 9, our theory’s conservative estimate of $K$ aligns with the required repetitions needed to achieve a robustness close to (but smaller than) $\rho=0.05$. Thus, the theory produces estimates that are geared towards more aggressive attacks.
@@
To test the impact of the non-random attack, we’ll next consider graphs with preferential attachment, generated using the Barabási-Albert (BA) model. Unlike the case of Erdős-Rényi (ER) graphs, BA graphs exhibit power law behavior in the vertex degree distributions making them more vulnerable to the non-random component of the attack (removing edges connected to the vertex with the highest eigenvector centrality). As this power law behavior is more pronounced in larger graphs, we’ll consider the case $N=1000$ and compare the effectiveness of the method with those of the ER model shown in Fig. 6a. The empirical distribution obtained by simulating graphs randomly using the BA model shifts to the right when compared to ER model as shown in Fig. 10, illustrating the need for a larger $K$ for graphs that have power law degree distributions.
(a) Erdős-Rényi graph
(b) Barabási-Albert graph
Figure 10: $\widehat{f}_{h,M}$ for (a) Erdős-Rényi (ER) versus (b) Barabási-Albert (BA) graphs. The settings of the experiment are the same as those used in Fig. 6. The PDFs associated with the BA graph are flatter and shifted to the right illustrating that BA graphs are more susceptible to attacks that target the network topology like the one considered here.
Finally, Fig. 11 shows the ECDF in the BA case for $N=1000$ and $K=10$. We only needed 10 repetitions for the distribution to converge, well below the theoretical limit given by (41), and we see that the $\rho=0.05$ robustness level was achieved with 99% probability. However, upon comparing this with Fig. 6a, we see a significant difference between ER graphs and BA graphs. This suggests that a more nuanced approach, which also considers graph topology could be more suitable for scale-free graphs.
Figure 11: $\widehat{F}_{M}$ for $N=1000$, $\nu=0.01$, $K=10$, and graphs generated with the BA model.
@@ VI Conclusion
For attacks that target specific edges (e.g., Xu et al. (2019)), the defender-added perturbation $\vec{\zeta}$ becomes necessary. It serves the role of concealing the graph structure from potential attackers. The bound on the number of repetitions derived above tends to be robust to more aggressive attacks. This is understandable as it aims to be robust to attacks of unknown nature. For scale-free graphs, we observe that larger numbers of repetitions are needed to achieve accurate decoding with very high probability (99%). This suggests that extending the present approach to incorporate elements of network structure could be an effective error correction mechanism for scale-free networks.